OpenStudy (anonymous):

Graph the function y=log(x+1)+1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

if you start with your base function \[f(x)=\log (x)\] what transformations do you apply to get it in the form\[y=\log(x+1) +1\]

OpenStudy (anonymous):

okay then what else do you do im so confused

OpenStudy (anonymous):

\[f(x+1)\] is a transformation 1 unit to the left\[f(x-1)\] is a transformation 1 unit to the right\[f(x)+1\] is a transformation 1 unit up\[f(x)-1\] is a transformation 1 unit down

OpenStudy (anonymous):

which of these has been applied to \[f(x)=\log(x)\]??

OpenStudy (anonymous):

 this is what the graph of y=log(x) looks like without any transformations